OpenStudy (anonymous):

graph the equation using the slope and y intercept; x-y=-10

OpenStudy (superdavesuper):

Y intercept is at x=0 so 0 - y = -10 i.e. y=10. Slope of the equation is given by rearranging equation to y = x + 1 i.e. slope=1 You can then graph the equation.